# Uncertainty Quantification in Retrieval Augmented Question Answering

### v2

@@ 1 Introduction
Retrieval augmented QA architectures have proven beneficial in increasing LLM performance on tail knowledge (Izacard et al., 2024; Mallen et al., 2023), reducing hallucinations in the generated answers (Shuster et al., 2021), and even improving model calibration (Jiang et al., 2021). However, there are various ways in which retrieval augmented QA can go wrong at inference time. The set of retrieved passages is far from perfect (Sciavolino et al., 2021; Yoran et al., 2024; Kasai et al., 2024) containing irrelevant or misleading evidence, the model might be under-trained to read certain passages and reason over these and the question (Izacard et al., 2024; Liu et al., 2024b), and the question can be ambiguous or unanswerable (Kasai et al., 2024). Ultimately, QA models may not follow the provided passages (Xie et al., 2024; Joren et al., 2025). When faced with uncertainty, QA models should ideally acknowledge it (e.g., by communicating it) rather than risk an incorrect response.
A good deal of previous work has focused on quantifying answer uncertainty in the context of closed-book QA tasks, where the answer is predicted based on a question and the model’s encoded knowledge. Sampling-based methods rely on output discrepancies among multiple predictors of the same input (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017). They measure diversity on a set of answers (Kuhn et al., 2023; Chen & Mueller, 2024) sampled via temperature scaling (Guo et al., 2017), with larger variance indicating higher uncertainty. LLM-based methods rely on the QA model’s own judgment of uncertainty (Kadavath et al., 2022; Lin et al., 2022; Tian et al., 2023). Through prompting, the model is encouraged to express its uncertainty (e.g., 0.5 or ‘almost certain’), either alongside the predicted answer (Lin et al., 2022; Tian et al., 2023) or after generating it (Kadavath et al., 2022; Tian et al., 2023). None of these approaches has been applied in the context of retrieval augmented QA.
@@ 2 Related Work
Several methods have been proposed to predict answer uncertainty in QA; however, none of them has analysed uncertainty in retrieval augmented QA models. Many existing approaches rely on the assumption that output variation is an expression of model uncertainty (Kuhn et al., 2023; Farquhar et al., 2024; Chen & Mueller, 2024). For example, Kuhn et al. (2023) first cluster answers with similar meaning (in a sample) via natural language inference before computing entropy while Chen & Mueller (2024) focus on black-box models; they also compute similarities in the set of answers but associate them with a model self-judgement of confidence. These approaches are expensive to run at inference time for a production QA system, they require several inference steps in addition to performing similarity computations which can become more complex with longer answers (Zhang et al., 2024b). Hou et al. (2024) focus on quantifying aleatoric uncertainty (i.e., uncertainty in the data) caused by ambiguous questions, an approach which could be combined with ours.
@@ 3 Modeling Answer Uncertainty
$P(y|x,R;\mathcal{M})=\prod_{t=1}^{|y|}P(y_{t}|y_{1..t-1},x,R;\mathcal{M}).$ (1)
@@ 3.1 Passage Utility Prediction
Accuracy, denoted as $a(y)$, indicates whether the generated answer $y$ is correct, while entailment, denoted as $e(y)$, measures the degree to which $p$ supports $y$. Accuracy is determined by an evaluator $A$, and entailment is assessed using a Natural Language Inference (NLI) classifier model $E$. The combined passage utility $\upsilon_{\mathcal{M}}$ ranges between 0 and 1, given that $a$ takes values in $\{0,1\}$ and $e$ falls within the $[0,1]$ interval.
We train a lightweight neural model on dataset $D_{\mathcal{M}}=\{(x,p,\upsilon_{\mathcal{M}})\}$ to predict passage utility scores, $\{x,p\}\mapsto\upsilon_{\mathcal{M}}(\{x,p\})$, We construct $D$ by pairing input questions $x$ and passages $p$ with utility scores $\upsilon_{\mathcal{M}}$ which we obtain after running $\mathcal{M}$ on examples ${(x,p)}$ and computing observed answer accuracy and entailment scores from the QA model $\mathcal{M}$. We retrieve $|R|>1$ passages per question to ensure a diverse range of usefulness and create training instances $\{(x,p_{i},\upsilon_{i})\,|\,p_{i}\in R\}$ under model $\mathcal{M}$. We leverage this data to train the passage utility predictor with a contrastive learning scheme. Specifically, if two passages $p_{i}$ and $p_{j}$ belong to $R$ and $p_{i}$ is more useful than $p_{j}$ for answering question $x$, the predicted utility score $\upsilon_{i}$ should be higher than $\upsilon_{j}$ by margin $m$, ensuring that $p_{i}$ is ranked above $p_{j}$. We train the utility predictor with the following ranking objective:
$\mathcal{L}_{rank}=\sum\limits_{((x,p_{i}),(x,p_{j}))\in R\times R,i\neq j}\max(0,-z(\upsilon_{i}-\upsilon_{j})+m)),$ (3)
where $z$ controls the gold order between $p_{i}$ and $p_{j}$ (e.g., if $z=1$, $p_{i}$ has higher utility, and conversely $z=-1$ indicates the opposite ordering) and $m$ is a hyper-parameter.
@@
$\mathcal{L}_{BCE}=\sum_{(x,p)\in\{(x,p_{i}),(x,p_{j})\}}[a\times\log(p(x,p))+(1-a)\times\log(1-p(x,p))],$ (4)
where $p(x,p)=\operatorname{sigmoid}(\upsilon)$ and $a$ is the target accuracy label under model $\mathcal{M}$ taking values in the set $\{0,1\}$. We train the utility predictor with the following combined objective:
$\mathcal{L}=\mathcal{L}_{rank}+\lambda\,\mathcal{L}_{BCE},$ (5)
where $\lambda$ is a hyper-parameter.
@@ 4 Experimental Setup
We evaluate our approach to predicting answer uncertainty on short-form question answering tasks. Specifically, on the following six datasets: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), WebQuestions (Berant et al., 2013), SQuAD (Rajpurkar et al., 2016), and PopQA (Mallen et al., 2023). We also evaluate on RefuNQ (Liu et al., 2024a), a dataset with unanswerable questions about non-existing entities. In Appendix E.1, we describe each dataset, provide example questions, and make available details about the splits used in our experiments which follow Lee et al. (2019).
QA Models
@@
We consider backbone instruction fine-tuned LLMs from different families of similar size. These are Llama-3.1-8B-Instruct (AI@Meta, 2024), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Gemma2-9B-it (Riviere et al., 2024). We also assess answer uncertainty quantification performance on QA models of the same family but with different sizes. To this end, we additionally evaluate on Gemma2-27B-it. For all QA models, we use a simple prompt including the retrieved passages and the question in the context; the prompt is shown in Table 9 in the Appendix. The QA models’ answer is the most likely answer generated with greedy sampling at temperature equal to 0. Following previous work on retrieval augmented QA, we use Contriever-MSMARCO (Izacard et al., 2022) as our external retriever (Asai et al., 2024) and the target QA models are prompted with $|R|=5$ passages (Yu et al., 2023; Asai et al., 2024; Xu et al., 2024). In Appendix E.2, we provide more details about inference and passage retrieval.
Accuracy Evaluation
@@
A precise metric for measuring accuracy is key when evaluating the quality of uncertainty estimation. Token overlap metrics are imprecise and can over- or under-estimate accuracy (e.g., 5 will not match five). Thus, we rely on a LLM-based accuracy evaluator to create training data for the Passage Utility predictor (i.e., $A$ in Section 3.1) and to evaluate retrieval augmented QA performance. We use Qwen2-72B-Instruct (Yang et al., 2024) and the prompt proposed by Sun et al. (2024) to obtain accuracy judgments. Details about the LLM evaluator can be found in Appendix E.2.
Evaluation of Uncertainty Estimation
@@
To assess the quality of answer uncertainty prediction, we follow Farquhar et al. (2024) and report the Area Under the Receiver Operator Curve (AUROC) on detecting incorrect answers (i.e., answers with high uncertainty). In Appendix F.2, we also report the area under the ‘rejection accuracy’ curve (AURAC) which captures the accuracy a model would have if it refused to answer questions with highest uncertainty. Rejection accuracy is essentially the model’s accuracy on the remaining questions. In the main results section, we focus on selective answering performance when models answer 80% of the least uncertain questions versus when always answering. We provide implementation details in Appendix E.2.
Training of Passage Utility Predictor
@@
We train a Passage Utility predictor per QA model and QA task. For each task, we create set $D_{\mathcal{M}}=\{(x,p,\upsilon_{\mathcal{M}})\}$ to train and evaluate a Passage Utility predictor for QA model $\mathcal{M}$. We use the training (and development) questions available for each QA task, considering the top five retrieved passages for each question (i.e., $|R|=5$). Note that $|R|$ a hyper-parameter and other values would be also possible. Larger sizes of $|R|$ would yield more training data, since the Utility predictor takes individual passages (together with the question) as input, The target QA model $\mathcal{M}$ is first prompted with passage $p\in R$ and question $x$ to generate answer $y$. Then, we annotate passages $p$ with a utility score computed with the accuracy evaluator $A$ and entailment judge $E$ on generated answer $y$ (Section 3.1). We use an ALBERT-xlarge model (Lan et al., 2020) optimized on MNLI (Williams et al., 2018) and the VitaminC dataset (Schuster et al., 2021). We report further details about the Passage Utility predictor training in Appendix E.2.
@@
There exist several uncertainty estimation methods which we group in two categories based on whether they require access to logits or simply model outputs (see Fadeeva et al. 2023 for additional methods). We choose the highest scoring ones to compare with here and include additional results in Appendix F.2 for completeness.
@@
LLM-based Measures. We compare with p(true) which uses the same LLM-based target QA model to assess whether the answers it produces are correct (Kadavath et al., 2022). We follow the p(true) variant used in previous work (Kuhn et al., 2023). The QA model is prompted with the question and a set of candidate answers (consisting of the most likely answer and a sample of $N$ answers) and is instructed to respond whether the most likely answer is true or false (i.e., correct/incorrect). The score produced by this approach is the probability of model $\mathcal{M}$ generating the token True. This method needs several in-context examples to work well; following Kuhn et al. (2023), we use 20 in-context examples. Note that since our backbone LLMs are recent models with a large context (unlike Kuhn et al. 2023), all 20 examples are fed in the context making p(true) an expensive but very strong approach. In the context of retrieval augmented QA, we modify p(true) to include in the prompt the set of retrieved passages for the question of interest. We provide the prompt used by p(true) in Appendix E.3. Note that p(true) can be considered as a specialized powerful entailment model and thus we do not include entailment based methods relying on off-the-shelf NLI models which have been shown to perform poorly (Yoran et al., 2024).
[…]
@@ 5 Results and Analysis
Table 1 summarizes our uncertainty estimation results (test set) with four QA models (Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3) across six QA datasets (results on the development set are included in Appendix F.2). We boldface the highest AUROC value for each QA model and dataset pair and mark with  the next best value that is significantly different from it at $p<0.05$. We use the paired De Long test (DeLong et al., 1988) to compute whether two AUROC values are significantly different. [^2] We use the library in https://github.com/Brritany/MLstatkit to compute significance scores.
@@
Utility Ranker 0.77 0.82 0.72 0.83 0.87 0.81 0.80 0.74 0.83 0.68 0.82 0.85 0.80 0.79
Table 1: AUROC values for QA models based on Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3 on Natural Questions (NQ), TriviaQA (TQA), WebQuestions (WebQ), SQuAD, PopQA, and RefuNQ test sets. The best values (per model and dataset) are highlighted in bold; we also mark with next best values which are significantly different using the paired De Long test ($p<0.05$). For example, on TQA with Gemma2-9B, p(true), the second best performing is significantly different from the Utility Ranker which performs best and by extension models with lesser values than p(true) are also significantly different.
@@
7009k13k0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8PPL (27B)p(true) (27B)SE (27B)PU (27B)PPL (9B)p(true) (9B)SE (9B)PU (9B)Number of Inference TokensAverage AUROCGemma2-9BGemma2-27B
@@
G9BG27BL8BM7B0.60.60.60.60.650.650.650.650.70.70.70.70.750.750.750.75Model variantsAverage QA AccuracyPPLRegular EntropyPassage Utilityp(true)Semantic Entropy
@@
Model uncertainty can be used to decide whether to provide an answer to a question or not. Figure 3 shows average accuracy when the target QA models choose to answer 80% of the cases they are most confident about. For comparison, we also show QA accuracy when always answering, i.e., black bold dots. All uncertainty quantification approaches improve model accuracy. Across different LLM QA backbones, Passage Utility performs on par with or better than more expensive uncertainty estimation approaches such as p(true). For instance, when looking at selective performance according to Passage Utility, the biggest Gemma2-27B model improves by +9 points (0.74 vs 0.65). In Appendix F.2 we report the full set of selective accuracies at different thresholds.
@@
Passage Utility generalizes to complex reasoning questions.
We assess the generalization of our approach on multi-hop questions, i.e., questions that require reasoning over pieces of information distributed across more than one passage (Yang et al., 2018; Pal et al., 2022). We evaluate on the widely used HotPotQA dataset (Yang et al., 2018), we use the splits as provided by Trivedi et al. (2023), and the same retrieval settings as defined above (Section 4). Table 2 shows AUROC values for all uncertainty estimation methods.
Across models, the performance of our approach is on par with p(true) and sequence entropy methods, and better than perplexity. Manual inspection of 100 examples from the development set with Gemma2-9B as the target QA model, reveals two major trends. Firstly, QA models can perform the multi-hop task by relying on memorized knowledge and the provided passages. In 48 cases, the QA model can correctly answer with several of the passages in $R$ when prompted individually; the model also correctly answers when prompted with $R$. Secondly, the retrieved passages tend to be of low quality. For 29 examples, the model cannot answer neither when prompted with individual passages nor with the entire set $R$. Our approach works well in these two cases with few false positives (negatives). Beyond the assessment of answer uncertainty estimation, this evaluation shows that the idealized multi-hop behavior studied with synthetic passages (Yang et al., 2018) occurs to a lesser extent in current LLMs in the context of open book QA where it is more challenging to retrieve relevant passages in the first place. Note that multi-hop QA often requires more complex QA pipelines such as interactive retrieval (Trivedi et al., 2023).
5310.50.50.50.50.60.60.60.60.70.70.70.7Number of Retrieved PassagesAverage AccuracyRAGRAG+PURAG+PPL

### v3

@@ 1 Introduction
Retrieval augmented QA architectures have proven beneficial in increasing LLM performance on tail knowledge (Izacard et al., 2024; Mallen et al., 2023), reducing hallucinations in the generated answers (Shuster et al., 2021), and even improving model calibration (Jiang et al., 2021). However, there are various ways in which retrieval augmented QA can go wrong. The set of retrieved passages is far from perfect (Sciavolino et al., 2021; Yoran et al., 2024; Kasai et al., 2024) containing irrelevant, incomplete, or misleading evidence, the model might be under-trained to read certain passages and reason over these and the question (Izacard et al., 2024; Liu et al., 2024b; Sun et al., 2025), and the question can be ambiguous or unanswerable (Kasai et al., 2024). Ultimately, QA models may not follow the provided passages (Xie et al., 2024; Joren et al., 2025). When faced with uncertainty, QA models should ideally acknowledge it (e.g., by communicating it) rather than risk an incorrect response.
A good deal of previous work has focused on quantifying answer uncertainty in the context of closed-book QA tasks, where the answer is predicted based on a question and the model’s encoded knowledge. Sampling-based methods rely on output discrepancies among multiple predictors of the same input (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017). They measure diversity on a set of answers (Kuhn et al., 2023; Chen & Mueller, 2024) sampled via temperature scaling (Guo et al., 2017), with larger variance indicating higher uncertainty. LLM-based methods rely on the QA model’s own judgment of uncertainty (Kadavath et al., 2022; Lin et al., 2022; Tian et al., 2023). Through prompting, the model is encouraged to express its uncertainty (e.g., 0.5 or ‘almost certain’), either alongside the predicted answer (Lin et al., 2022; Tian et al., 2023) or after generating it (Kadavath et al., 2022; Tian et al., 2023). None of these approaches has been applied in the context of retrieval augmented QA (Shorinwa et al., 2025).
@@ 2 Related Work
Several methods have been proposed to predict answer uncertainty in the context of closed-book QA; however, none of them has analysed uncertainty in retrieval augmented QA models. Many of these methods rely on the assumption that variation in the model’s output is an expression of its uncertainty (Kuhn et al., 2023; Farquhar et al., 2024; Chen & Mueller, 2024). For example, some approaches (Kuhn et al., 2023; Farquhar et al., 2024) first cluster answers with similar meaning (in a sample) via Natural Language Inference (NLI) before computing entropy while other work (Chen & Mueller, 2024) focuses on black-box models; they also compute similarities in the set of answers but associate them with the model’s self-judgement of confidence. In a recent study, Soudani et al. (2025) show that, in the context of retrieval augmented QA, these methods tend to exhibit overconfidence and sensitivity to the input context. Indeed, our experiments corroborate their analysis. We find that sampling-based methods exhibit less variation in a retrieval setting, and do not offer a significant advantage over methods such as perplexity or LLM prompting for self-assessment (Kadavath et al., 2022), unlike in closed-book settings.
Sequence entropy methods (Kuhn et al., 2023; Farquhar et al., 2024) focus on detecting incorrect answers stemming from arbitrary fluctuations in model outputs (referred to as confabulations). Our approach extends to additional error sources, including incorrect training data or misleading evidence. Hou et al. (2024) focus on quantifying aleatoric uncertainty (i.e., uncertainty in the data) caused by ambiguous questions, an approach that could be combined with ours. Sampling-based methods are expensive to run at inference time for a production QA system; they require several inference steps in addition to performing similarity computations, which can become more complex with longer answers (Zhang et al., 2024b). In contrast, our approach is light-weight at inference time, but requires training data with observations pertaining to the accuracy of the target model. Our approach optionally uses an NLI model at training time, depending on the chosen objective. In Section 6, we further summarise key differences between our approach and existing uncertainty estimation methods in the context of retrieval augmented QA.
@@ 3 Modeling Answer Uncertainty
$\hskip-2.15277ptP(y|x,R;\mathcal{M})=\prod_{t=1}^{|y|}P(y_{t}|y_{1..t-1},x,R;\mathcal{M}).$ (1)
@@ 3.1 Passage Utility Prediction
Accuracy, denoted as $a(y)$, indicates whether the generated answer $y$ is correct, while entailment, denoted as $e(y)$, measures the degree to which $p$ supports $y$. Accuracy is determined by an evaluator $A$, and entailment is assessed using an NLI model $E$. The combined passage utility $\upsilon_{\mathcal{M}}$ ranges between 0 and 1, given that $a$ takes values in $\{0,1\}$ and $e$ falls within the $[0,1]$ interval. The intuition behind this composite term $v_{\mathcal{M}}$ is to encourage a ranking of passages that spans a broad spectrum: from highly useful ones (producing correct answers with strong entailment), through more ambiguous cases (e.g., correct answers with weak entailment or incorrect answers with strong entailment), to clearly unhelpful passages (producing incorrect answers with weak entailment). [^2] Note that for the pairwise ranking loss, using the sum $\upsilon_{\mathcal{M}}=a(y)+e(y)$ would be equivalent. In Appendix A, we include ablation experiments with different implementations of the Passage Utility score $\upsilon_{\mathcal{M}}$.
We train a lightweight neural model on dataset $D_{\mathcal{M}}=\{(x,p,\upsilon_{\mathcal{M}})\}$ to predict passage utility scores, $\{x,p\}\mapsto\upsilon_{\mathcal{M}}(\{x,p\})$, We construct $D$ by pairing input questions $x$ and passages $p$ with utility scores $\upsilon_{\mathcal{M}}$ which we obtain after running $\mathcal{M}$ on examples (x,p){(x,p)} and computing observed answer accuracy and entailment scores from the QA model $\mathcal{M}$. We retrieve $|R|>1$ passages per question to ensure a diverse range of usefulness and create training instances $\{(x,p_{i},\upsilon_{i})\,|\,p_{i}\in R\}$ under model $\mathcal{M}$. We leverage this data to train the passage utility predictor with a contrastive learning scheme. Specifically, if two passages $p_{i}$ and $p_{j}$ belong to $R$ and $p_{i}$ is more useful than $p_{j}$ for answering question $x$, the predicted utility score $\upsilon_{i}$ should be higher than $\upsilon_{j}$ by margin $m$, ensuring that $p_{i}$ is ranked above $p_{j}$. We train the utility predictor with the following ranking objective:
$\mathcal{L}_{rank}=\hskip-28.45274pt\sum\limits_{((x,p_{i}),(x,p_{j}))\in R\times R,i\neq j}\hskip-28.45274pt\max(0,-z(\upsilon_{i}-\upsilon_{j})+m)),$ (3)
where $z$ controls the gold order between $p_{i}$ and $p_{j}$ (e.g., if $z=1$, $p_{i}$ has higher utility, and conversely z=1z=-1 indicates the opposite ordering) and $m$ is a hyper-parameter.
@@
$\mathcal{L}_{BCE}=\sum\limits_{((x,p_{i}),(x,p_{j}))\in R\times R,i\neq j}[a\times\log(p(x,p_{i}))+(1-a)\times\log(1-p(x,p_{i}))]$ (4)
$+[a\times\log(p(x,p_{j}))+(1-a)\times\log(1-p(x,p_{j}))],$
where $p(x,p_{i})=\operatorname{sigmoid}(\upsilon_{i})$ and $a$ is the target accuracy label $a(y)$ under model $\mathcal{M}$ taking values in the set $\{0,1\}$. We train the utility predictor with the following combined objective where $\lambda$ is a hyper-parameter to adjust the $\mathcal{L}_{BCE}$ penalty:
$\mathcal{L}=\mathcal{L}_{rank}+\lambda\,\mathcal{L}_{BCE}.$ (5)
@@ 4.1 QA Tasks and Models
We evaluate our approach to predicting answer uncertainty on short-form question answering tasks. Specifically, on the following six datasets: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), WebQuestions (Berant et al., 2013), SQuAD (Rajpurkar et al., 2016), and PopQA (Mallen et al., 2023). We also evaluate on RefuNQ (Liu et al., 2024a), a dataset with unanswerable questions about non-existing entities. In Appendix F.1, we describe each dataset, provide example questions, and make available details about the splits used in our experiments which follow Lee et al. (2019).
@@
We consider backbone instruction fine-tuned LLMs from different families of similar size. These are Llama-3.1-8B-Instruct (AI@Meta, 2024), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Gemma2-9B-it (Riviere et al., 2024). We also assess answer uncertainty quantification performance on QA models of the same family but with different sizes. To this end, we additionally evaluate on Gemma2-27B-it. For all QA models, we use a simple prompt including the retrieved passages and the question in the context; the prompt is shown in Table 14 in the Appendix. The QA models’ answer is the most likely answer generated with greedy sampling at temperature equal to 0. Following previous work on retrieval augmented QA, we use Contriever-MSMARCO (Izacard et al., 2022) as our external retriever (Asai et al., 2024) and the target QA models are prompted with $|R|=5$ passages (Yu et al., 2023; Asai et al., 2024; Xu et al., 2024). In Appendix F.2, we provide more details about inference and passage retrieval.
4.2 Evaluation
QA Accuracy
@@ 4.2 Evaluation
A precise metric for measuring accuracy is key when evaluating the quality of uncertainty estimation. Token overlap metrics are imprecise and can over- or under-estimate accuracy (e.g., 5 will not match five). Thus, we rely on a LLM-based accuracy evaluator to create training data for the Passage Utility predictor (i.e., $A$ in Section 3.1) and to evaluate retrieval augmented QA performance. We use Qwen2-72B-Instruct (Yang et al., 2024) and the prompt proposed by Sun et al. (2024) to obtain accuracy judgments. Details about the LLM evaluator can be found in Appendix F.2.
Uncertainty Estimation
@@
To assess the quality of answer uncertainty prediction, we follow Farquhar et al. (2024) and report the Area Under the Receiver Operator Curve (AUROC) on detecting incorrect answers (i.e., answers with high uncertainty). In Appendix G.3, we also report the area under the ‘rejection accuracy’ curve (AURAC) which captures the accuracy a model would have if it refused to answer questions with highest uncertainty. Rejection accuracy is essentially the model’s accuracy on the remaining questions. In the main results section, we focus on selective answering performance when models answer 80% of the least uncertain questions versus when always answering. We provide implementation details in Appendix F.2.
4.3 Methods
Passage Utility Predictor
@@ 4.3 Methods
We train a Passage Utility predictor per QA model and QA task. For each task, we curate dataset $D_{\mathcal{M}}=\{(x,p,\upsilon_{\mathcal{M}})\}$ to train and evaluate a Passage Utility predictor for QA model $\mathcal{M}$. We use the training (and development) questions available for each QA task, considering the top five retrieved passages for each question (i.e., $|R|=5$). Note that $|R|$ is a hyper-parameter and other values would be also possible. Larger sizes of $|R|$ would yield more training data, since the Utility predictor takes individual passages (together with the question) as input. The target QA model $\mathcal{M}$ is first prompted with passage $p\in R$ and question $x$ to generate answer $y$. Then, we annotate passages $p$ with a utility score computed with the accuracy evaluator $A$ and entailment judge $E$ on generated answer $y$ (Section 3.1). We use an ALBERT-xlarge model (Lan et al., 2020) optimized on MNLI (Williams et al., 2018) and the VitaminC dataset (Schuster et al., 2021). We provide more details about the curated datasets and training of the Passage Utility predictor training in Appendix F.2.
@@
There exist several uncertainty estimation methods which we group in two categories based on whether they require access to logits or simply model outputs (see Fadeeva et al. 2023 for additional methods). We choose the highest scoring ones to compare with here and include additional results in Appendix G.3 for completeness.
@@
LLM-based Measures. We compare with p(true) which uses the same LLM-based target QA model to assess whether the answers it produces are correct (Kadavath et al., 2022). We follow the p(true) variant used in previous work (Kuhn et al., 2023). The QA model is prompted with the question and a set of candidate answers (consisting of the most likely answer and a sample of $N$ answers) and is instructed to respond whether the most likely answer is true or false (i.e., correct/incorrect). The score produced by this approach is the probability of model $\mathcal{M}$ generating the token True. This method needs several in-context examples to work well; following Kuhn et al. (2023), we use 20 in-context examples. Note that since our backbone LLMs are recent models with a large context (unlike Kuhn et al. 2023), all 20 examples are fed in the context making p(true) an expensive but very strong approach. In the context of retrieval augmented QA, we modify p(true) to include in the prompt the set of retrieved passages for the question of interest. We provide the prompt used by p(true) in Appendix F.3. Note that p(true) can be considered as a specialized powerful entailment model and thus we do not include entailment based methods relying on off-the-shelf NLI models which have been shown to perform poorly (Yoran et al., 2024).
For approaches that require sampling, we follow previous work (Farquhar et al., 2024) and take $N=10$ samples, which we generate with multinomial sampling. We set the sampling temperature to 1, with nucleus sampling ($P=0.9$; Holtzman et al. 2020) and topK-K sampling ($K=50$; Fan et al. 2018), and use a different random seed to draw each sample. We provide further details about inference in Appendix F.2 and report inference costs for each approach in Appendix B.
@@ 5 Results and Analysis
Table 1 summarizes our uncertainty estimation results (test set) with four QA models (Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3) across six QA datasets (results on the development set are included in Appendix G.3). We boldface the highest AUROC value for each QA model and dataset pair and mark with  the next best value that is significantly different from it at $p<0.05$. We use the paired De Long test (DeLong et al., 1988) to compute whether two AUROC values are significantly different. [^3] We use the library in https://github.com/Brritany/MLstatkit to compute significance scores.
@@
Passage Utility 0.77 0.82 0.72 0.83 0.87 0.81 0.80 0.74 0.83 0.68 0.82 0.85 0.80 0.79
Table 1: AUROC values for QA models based on Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3 on Natural Questions (NQ), TriviaQA (TQA), WebQuestions (WebQ), SQuAD, PopQA, and RefuNQ test sets. The best values (per model and dataset) are highlighted in bold; we also mark with next best values which are significantly different using the paired De Long test ($p<0.05$). For example, on TQA with Gemma2-9B, p(true), the second best performing is significantly different from the Passage Utility which performs best and by extension models with lesser values than p(true) are also significantly different.
@@
7009k13k0.50.50.60.60.70.70.80.8PPLp(true)SEPUPPLp(true)SEPUNumber of Inference TokensAverage AUROCGemma2-9BGemma2-27B
@@
G9BG27BL8BM7B0.60.60.650.650.70.70.750.75Model variantsAverage QA AccuracyPPLRegular EntropyPassage Utilityp(true)Semantic Entropy
@@
Model uncertainty can be used to decide whether to provide an answer to a question or not. Figure 3 shows average accuracy when the target QA models choose to answer 80% of the cases they are most confident about. For comparison, we also show QA accuracy when always answering, i.e., black bold dots. All uncertainty quantification approaches improve model accuracy. Across different LLM QA backbones, Passage Utility performs on par with or better than more expensive uncertainty estimation approaches such as p(true). For instance, when looking at selective performance according to Passage Utility, the biggest Gemma2-27B model improves by +9 points (0.74 vs 0.65). In Appendix G.3 we report the full set of selective accuracies at different thresholds.
@@
Passage Utility struggles with multi-hop questions.
Our approach estimates uncertainty by predicting individual passage utility and selecting the maximum utility score from a set of retrieved passages. A potential limitation of this method emerges in multi-hop questions that require evidence from multiple passages (Yang et al., 2018; Pal et al., 2022). Specifically, the QA model is unable to answer correctly when prompted with any individual passage (resulting in uniformly low utility scores), yet succeeds when given the complete set of passages (in which case the maximum utility score over individual passages fails to reflect this outcome). To quantify this limitation, we evaluate our approach on the widely used HotPotQA dataset (Yang et al., 2018), using the splits as provided by Trivedi et al. (2023), and the same retrieval settings as defined above (Section 4). Table 2 shows AUROC values for all uncertainty estimation methods. In Appendix E, we provide an analysis on a small set of multi-hop questions making use of synthetic HotPotQA data.
Across models, our approach is on par with p(true) and sequence entropy, and better than perplexity (Table 2). Manual inspection of 100 examples from the development set with Gemma2-9B as the target QA model, reveals two major trends. Firstly, the QA model frequently (48 cases) manages to correctly answer multi-hop questions using only one of the required ’hop’ passages, often not needing the entire set. This observation aligns with recent findings in Joren et al. (2025). Secondly, in numerous instances (29 cases), the retrieved passages did not contain any useful evidence, leading the model to answer incorrectly even when prompted with the full set. This underscores the inherent difficulty of effective retrieval for complex questions. Many studies (Jeong et al., 2024; Trivedi et al., 2023; Lin et al., 2025) tackle this challenge through sophisticated QA pipelines. Notably, certain approaches decompose complex questions into sub-questions that can be answered independently. In such multi-hop pipelines, our approach could be naturally applied at the sub-question level.
5310.50.50.60.60.70.7Number of Retrieved PassagesAverage AccuracyRAGRAG+PURAG+PPL
@@
6 Discussion
Key Properties and Usability Scenarios
Each uncertainty quantification approach comes with its own advantages and limitations. This entails that the choice of a specific method depends on criteria like available resources, desired latency, and the necessary level of control and trustworthiness for the QA system. For example, in high-stakes applications, a method that favours higher rates of false positives (thereby allowing human intervention) and reduces the chance of overconfident false negatives would be better, even if it requires training data and regular updates.
In Table 3, we summarise existing features (and limitations) of various uncertainty estimation approaches as we compare them to our work. The first column reports latency, as previously discussed in Figure 2. Retrieval augmented QA models can err as a result of being exposed to erroneous sources such as misleading passages (Xie et al., 2024) or inaccurate training data (Vu et al., 2024). Information-theoretic methods are not equipped with an explicit mechanism to deal with these cases (Farquhar et al., 2024; Soudani et al., 2025). Moreover, these methods are also known to suffer from over-confidence (Simhi et al., 2025; Soudani et al., 2025; Sung et al., 2025). While p(true) may be able to detect these challenging cases to a certain extent, Passage Utility can be specifically trained to recognise them. In terms of supervision, both p(true) and Passage Utility require task-specific training examples, and the performance of both approaches deteriorates on out-of-distribution examples (Table 8).
Low latency No Training Data No Fine-tuning Recognises Erroneous Evidence Mitigates overconfidence
PPL
p(true)
Semantic Entropy
Passage Utility
Table 3: Categorization of the uncertainty estimation approaches studied in Section 5 according to different properties (table headers).
Training Data Requirements
Our approach requires question-answer pairs to curate a dataset with retrieved passages and utility scores for training. However, general or task specific training datasets could be generated semi-automatically Li & Zhang (2024); Wei et al. (2024). Moreover, in experiments we show that in some QA tasks such as WebQ or PopQA this training data can be relatively small, i.e., in the region of 2.5k or 10k respectively.
Fine-Tuning Requirements
Our approach, by design, requires fine-tuning to adapt to new QA tasks or models, as its core aim is to model the accuracy behavior of the target QA model. To enhance versatility across QA tasks, a unified training set encompassing diverse QA tasks could be compiled to train a single passage utility predictor. More practically, advanced training schemes promoting generalization, such as meta-learning with a varied set of QA tasks and QA model examples, could be employed to develop a single passage utility predictor.
In addition, while the utility predictor may necessitate recalibration for distinct QA models or tasks, its performance on out-of-distribution scenarios (Appendix D) establishes it as a robust warm-up base model. This allows for subsequent fine-tuning with a reduced number of target examples (Kamath et al., 2020; Zhang et al., 2021). Finally, as the utility predictor relies on a small model, the cost of fine-tuning in terms of resources and time is low.